Magnesium chloride, MgCl2, is a common roadway deicer salt. calculate the weight percent of the chloride in the salt. show your work.
Answer #1


Molar mass of MgCl2,
MM = 1*MM(Mg) + 2*MM(Cl)
= 1*24.31 + 2*35.45
= 95.21 g/mol


MM of MgCl2 = 95.21
Lets calculate the mass of element

mass of Cl = 2*MM(Cl)
= 70.9 g

Mass % of Cl = (70.9*100)/95.21
= 74.47 %
Answer: 74.47 %